# FLUKE: A Linguistically-Driven and Task-Agnostic Framework for Robustness Evaluation

### v2

@@ Abstract
We present FLUKE (framework for linguistically-driven and task-agnostic robustness evaluation), a framework for assessing model robustness through systematic minimal variations of test data. FLUKE introduces controlled variations across linguistic levels — from orthography to dialect and style — and leverages large language models (LLMs) with human validation to generate modifications. We demonstrate FLUKE’s utility by evaluating both fine-tuned models and LLMs across six diverse NLP tasks (four classification and two generation tasks), and reveal that (1) the impact of linguistic variations is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) LLMs still exhibit significant brittleness to certain linguistic variations, with reasoning LLMs surprisingly showing less robustness on some tasks compared to base models; (3) models are overall more brittle to natural, fluent modifications such as syntax or style changes (and especially to negation), compared to corruption-style tests such as letter flipping; (4) the ability of a model to use a linguistic feature in generation does not correlate to its robustness to this feature on downstream tasks. These findings highlight the importance of systematic robustness testing for understanding model behaviors.
@@
Yulia Otmakhova1∗    Hung Thinh Truong1∗    Rahmad Mahendra 2    Zenan Zhai 3 Rongxin Zhu 1,3    Daniel Beck 2    Jey Han Lau1 1The University of Melbourne    2RMIT University    3Oracle {y.otmakhova, thinh.truong}@unimelb.edu.au    rahmad.mahendra@student.rmit.edu.au zenanzhai@gmail.com    rongxin.zhu@unimelb.edu.au     daniel.beck@rmit.edu.au jeyhan.lau@unimelb.edu.auu
@@ 1 Introduction
To demonstrate the generalisability of FLUKE, we apply it to modify instances for four classification tasks of various complexity — coreference resolution, dialogue contradiction classification, named entity recognition (NER), and sentiment analysis — as well as two generation tasks — instruction-following evaluation (IFEval) and grade school math (GSM). We then test a number of models — including fine-tuned pre-trained models (PLMs: BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019) and T5 (Raffel et al., 2020)), base LLMs (GPT-4o, Llama 3.1, and Claude 3.5), as well as reasoning LLMs (GPT-5 and DeepSeek R1) — and examine their predictions in the original and modified instances. When measuring the overall robustness of the model, we account for two cases: where the label of the instance is supposed to stay the same after modification and where the label changes. Accordingly, we consider the model to be robust if its predictions either remain the same or change together with the label. We find that: (1) Robustness to modifications highly depends on the task. (2) Even though LLMs are more robust than PLMs, they are still brittle to many modifications, and for some tasks the brittleness is surprisingly more prominent in reasoning LLMs. (3) Natural, linguistically equivalent modifications are more damaging than traditional adversarial modifications such as random punctuation insertion; and (4) Models can be robust to linguistic features that they struggled to understand when generating, and brittle to features that are easy for them to employ. Thus, using only a handful of well-known tests may fail to reveal the brittleness of a model, highlighting the necessity of a comprehensive, task-agnostic testing framework like FLUKE. To summarise, our contributions are:
@@
  • We release our prompt templates, evaluation scripts, and validated modifications, to enable other researchers to test their models and tasks. [^3] https://anonymous.4open.science/r/fluke-BF6D We also publish a dashboard that allows to assess models performance by comparing their outputs on original and modified instances. [^4] https://fluke-viewer.pages.dev/
@@ 2 Related Work
Another way to assess models is through linguistic competence. A common approach is to use probing classifiers, tuned on datasets tailored to specific competences, such as morphology or syntax (Blevins et al., 2023; Mahowald et al., 2024). Holmes (Waldis et al., 2024) attempt to unify multiple datasets created to test linguistic competences. While probing for linguistic competence can provide insights about the model’s understanding of specific linguistic phenomena, it is important to note that model competence does not necessarily translate to task performance, as we show below in Section˜6.4.
@@ 3 FLUKE
FLUKE covers phenomena related to different levels of language structure, as well as some model biases. Simplified examples of each test are in Table 1.
@@ 4.1 Test set creation and validation
We generate modifications (Section˜3) for each task, randomly selecting samples from a test partition of the original dataset and prompting GPT-4o with instructions. We first create and evaluate a general prompt template for each modification, and then adapt it for each task (see Section˜A.1 for an example). We follow several steps to ensure the high quality of modifications. To improve their diversity, we stratify modifications across subtests (for example, different types of negation) and, where possible, randomly select a variable from a list (such as region and country for geographical bias or non-existing word for nonce semantic test) and pass it to the prompt. To ensure that modifications were applied minimally, i.e. the difference in performance can be attributed to a particular linguistic feature change rather than to a large number of changes, we calculate Levenshtein distance between original and modified samples, discarding those which changed more than expected. Finally, we employ two-stage human checks to ensure that modifications are valid. For classification tasks we ran this validation on Prolific [^7] https://www.prolific.com/, with thorough quality control and four annotators per sample (see Section˜A.2 for annotation details, inter-annotator agreement and quality statistics). During the first stage, we check if resulting texts have the required modification and are still fluent, rejecting samples that do not meet these criteria. We find that most tests have a high retain rate of 70-90%. During the second stage, we ask annotators to select the tasks’ label (such as positive or negative for Sentiment analysis) for modified samples, without knowing the labels for original samples, to check if the ground-truth label changed (e.g. after applying negation the sentiment can change from positive to negative). To make sure the task is solvable after modification, i.e. a label is not ambiguous or hard to choose, we add an option such as “Neutral” for Sentiment analysis or “Not sure” for other tasks, and reject samples where annotators chose it.
For generation tasks, to show how to apply the framework with minimal human intervention, we skip the first stage of human checks and only use tests with reliable modifications (over 70% retain rate). For the second stage, the authors of the paper validate modified samples in terms of whether the modification changes the label (for example, if it changes the premises in the math task, leading to a different answer) and whether the task is still solvable (for example, if there are modifications that prevent correct instruction following); see examples of valid and rejected modifications, as well as modifications that led to label change, in Table˜8 in Appendix. Such setup reduces the human effort, making the framework generalizable to new tasks, while still ensuring the validity of modifications.
@@ 4.2 Tested models and evaluation
We experiment with three smaller PLMs (BERT, GPT-2, and T5), three base LLMs (GPT-4o, Llama 3.1 and Claude 3.5), and two reasoning LLMs (GPT-5, DeepSeek R1). The pretrained models are fine-tuned on the training set with an additional task-specific classification layer on top (Section˜A.5); while the LLMs are prompted with instructions (zero-shot) (Table˜9).
@@
We follow the original metrics used for the task to define correctness. For Dialogue, Coreference, Sentiment, GSM the correctness metric is accuracy. For NER, we use per-sample entity-level mean F1 (Nakayama, 2018), while IFEval employs strict success (all constraints satisfied) as correctness.
@@ 5.1 Use Case: NER and Coreference tasks
Table 2: NER: Unrobustness (U, %) by model and modification. Claude stands for Claude 3.5, Llama for Llama 3.1, and DS for DeepSeek R1
@@
Table 3: Coreference: Unrobustness (U, %) by model and modification. Notation as in Table˜2
@@ 5.2 Use case: GSM and IFEval
We find that on the GSM task (Table˜4) models demonstrate major brittleness only to Negation tasks, where all of them fail to take into account the changed or even reversed logic of the premise, showing memorisation effects. Base LLMs such as Claude-3.5 and especially Llama 3.1 are also affected by context changes, such as failing to accomodate for changed names and currencies in Geographical bias test, or explaining the problem in informal style or Singlish dialect. Interestingly, reasoning LLMs are more robust to such changes.
@@
Table 4: GSM: Unrobustness (U, %) by model and modification
@@
Table 5: IFEval: Unrobustness (U, %) by model and modification
@@ 6.1 Model brittleness is task-specific
We call our framework task-agnostic; this, however, does not mean that all tests are meaningful for all the tasks. As the results above (and in Section˜A.6) show, the tests that actually “fired", or showed substantial drops or increases, are different for each task. Even test such as negation which is known to be difficult for models (Ravichander et al., 2022; Truong et al., 2022) has almost no effect on the NER task. On the other hand, seemingly “simple” tests (capitalization) that have a negligible effect on most of the tasks led to some dramatic drops on the NER and IFEval tasks. Thus, our framework is task-agnostic because it can be applied to reveal lack of robustness on a new task without having a preconceived (and potentially wrong) notion of what tests are meaningful for it.
@@ 6.2 LLMs are not always more robust than PLMs, and reasoning enhanced training does not always help
In classification tasks, where PLMs were finetuned on the original dataset, one would expect larger fluctuations in performance compared to LLMs, as modifications are likely to change the data distribution and break the identical distribution assumption that the fine-tuning process is based on. This holds for two out of four classification tasks we tested — Coreference resolution (Table˜3) and Sentiment Analysis (Table˜12 in Section˜A.6) — where PLMs are overall less robust to perturbations. However, on two other classification tasks — NER (Table˜2) and Dialogue contradiction detection (Table˜11 in Section˜A.6) — while base LLMs are overall more robust than PLMs, reasoning LLMs are as brittle as the latter. Moreover, while LLMs are on average (across all models and tests) better (see the “Average” row in Table˜14), they are sometimes markedly brittle to a particular test. For example, Claude-3.5 is substantially more affected by concept modifications than any PLMs on coreference resolution, or GPT-4o affected by grammatical role modifications on dialogue contradiction task.
Comparing between base and reasoning LLMs across tasks does not reveal a clear winner either. Reasoning seems to improve overall robustness on tasks such as GSM (Table˜4), Coreference (Table˜3) or Sentiment Analysys (Table˜12); however, on other tasks the reasoning models perform similar to base LLMs (IFEval, Table˜5) or even worse than them (NER, Table˜2, and Dialogue, Table˜11).
We also test if a reasoning model (GPT-5) is able to handle the modifications better if it is aware of them. To do that, we include additional context into the prompt where we alert the model that the sample was modified, and explicitly specify the type of modification (e.g. by telling it that the active voice was changed to passive). We found no consistent gain when we do this (see Appendix Table˜13).
@@ 6.3 Models are overall less robust to natural, linguistically valid modifications
Some of the work on robustness focuses on corrupting the text in adversarial way, such as shuffling characters or changing the case (Belinkov and Bisk, 2017; Li et al., 2018; Eger and Benz, 2020), while some includes natural, lingustically equivalent perturbations such as synonym replacement (Ren et al., 2019; Jin et al., 2020). Our framework combines both types of tests and allows to compare their efficiency for a particular task. We find that on average linguistically valid modifications reveal unrobustness more often than corrupting the text (i.e. such tests as Orthography) (see “Average” column in Table˜14). The exception here are the NER task where orthography is directly relevant to performance, and, surprisingly, IFEval, where the model often confused the instructions regarding capitalization with capitalization used in the instruction. Thus, the choice of the test type – adversarial corruptions or natural modifications – should also be motivated by the downstream task.
@@ 7 Conclusion
In this paper, we present FLUKE, a task-agnostic, linguistically-driven framework that introduces minimal modifications to existing datasets to create a benchmark for assessing robustness. We automate these modifications by prompting GPT-4o, followed by thorough manual validation to ensure data quality. To showcase the generalizability of this framework, we apply it to create modified instances across six tasks covering both classification and generation, and evaluate the performance of smaller fine-tuned models and large language models. Our findings reveal that: (1) the impact of modifications is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) while LLMs demonstrate better overall robustness compared to PLMs, they still exhibit significant brittleness, and on some tasks reasoning LLMs are less robust than base LLMs; (3) models are overall less robust to linguistically valid modifications than to adversarial corruption, but this again depends on the task; (4) the ability of a model to implement a linguistic feature during generation does not correlate with its robustness to that feature on downstream tasks. We hope that FLUKE provides an alternative approach for evaluating models, and that model developers would consider integrating these results into model cards (Mitchell et al., 2019) when releasing models.

### v3

@@ Abstract
We present FLUKE (framework for linguistically-driven and task-agnostic robustness evaluation), a framework for assessing model robustness through systematic minimal variations of test data. FLUKE introduces controlled variations across linguistic levels — from orthography to dialect and style — and leverages large language models (LLMs) with human validation to generate modifications. We demonstrate FLUKE’s utility by evaluating both fine-tuned models and LLMs across six diverse NLP tasks (four classification and two generation tasks), and reveal that (1) the impact of linguistic variations is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) LLMs still exhibit significant brittleness to certain linguistic variations, with reasoning LLMs surprisingly showing less robustness on some tasks compared to base models, and scaling improving robustness only for surface-level modifications; (3) models are overall more brittle to natural, fluent modifications such as syntax or style changes (and especially to negation), compared to corruption-style tests such as letter flipping; (4) the ability of a model to use a linguistic feature in generation does not correlate to its robustness to this feature on downstream tasks. These findings highlight the importance of systematic robustness testing for understanding model behaviors.
@@
Yulia Otmakhova1∗    Hung Thinh Truong1∗    Rahmad Mahendra 3 Zenan Zhai 2 Rongxin Zhu 2    Daniel Beck 4    Jey Han Lau1 1The University of Melbourne    2Oracle    3Universitas Indonesia    4RMIT University {y.otmakhova, thinh.truong}@unimelb.edu.au    rahmad.mahendra@cs.ui.ac.id zenanzhai@gmail.com    rongxin.zhu@unimelb.edu.au, daniel.beck@rmit.edu.au,     jeyhan.lau@unimelb.edu.au
@@ 1 Introduction
To demonstrate the generalisability of FLUKE, we apply it to modify instances for four classification tasks of various complexity — coreference resolution, dialogue contradiction detection, named entity recognition (NER), and sentiment analysis — as well as two generation tasks — instruction-following evaluation (IFEval) and mathematical reasoning (GSM). We then test a number of models — including fine-tuned pre-trained models (PLMs: BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019) and T5 (Raffel et al., 2020)), base LLMs (GPT-4o, Llama 3.1-405b, and Claude 3.5), as well as reasoning LLMs (GPT-5 and DeepSeek R1) — and examine their predictions in the original and modified instances. When measuring the overall robustness of the model, we account for two cases: where the label of the instance is supposed to stay the same after modification and where the label changes. Accordingly, we consider the model to be robust if its predictions either remain the same or change together with the label. We find that: (1) Robustness to modifications highly depends on the task. (2) Even though LLMs are more robust than PLMs, they are still brittle to many modifications, and for some tasks the brittleness is surprisingly more prominent in reasoning LLMs. (3) Scaling helps to improve robustness only to surface-level modifications. (4) Natural, linguistically equivalent modifications are more damaging than traditional adversarial modifications such as random punctuation insertion; and (5) Models can be robust to linguistic features that they struggled to understand when generating, and brittle to features that are easy for them to employ. Thus, using only a handful of well-known tests may fail to reveal the brittleness of a model, highlighting the necessity of a comprehensive, task-agnostic testing framework like FLUKE. To summarise, our contributions are:
@@
  • We release our prompt templates, evaluation scripts, and validated modifications, to enable other researchers to test their models and tasks. [^3] https://fluke-nlp.github.io/ We also publish a dashboard that allows to assess models performance by comparing their outputs on original and modified instances. [^4] https://fluke-viewer.pages.dev/
@@ 2 Related Work
Another way to assess models is through linguistic competence. A common approach is to use probing classifiers, tuned on datasets tailored to specific competences, such as morphology or syntax (Blevins et al., 2023; Mahowald et al., 2024). Holmes (Waldis et al., 2024) attempt to unify multiple datasets created to test linguistic competences. While probing for linguistic competence can provide insights about the model’s understanding of specific linguistic phenomena, it is important to note that model competence does not necessarily translate to task performance, as we show below in Section˜6.5.
@@ 3 FLUKE
FLUKE systematically covers phenomena related to different levels of language structure, as well as some model biases. Simplified examples of each test are in Table 1.
@@ 4 Task Evaluation
We apply modifications to test model robustness for six tasks, including four classification tasks (coreference resolution, dialogue contradiction detection, named entity recognition, sentiment analysis) and two generation tasks (instruction-following evaluation, IFEval, and mathematical reasoning, GSM).
@@ 4.1 Test set creation and validation
We generate modifications (Section˜3) for each task, randomly selecting samples from a test partition of the original dataset and prompting GPT-4o with instructions. We create and evaluate a general prompt template for each modification (see Section˜A.1 for an example), and then use the same prompt template for each task with minimal modifications (such as specifying the text type).
We follow several steps to ensure the high quality of modifications. To improve their diversity, we stratify modifications across subtests (for example, different types of negation) and, where possible, randomly select a variable from a list (such as region and country for geographical bias or non-existing word for nonce semantic test) and pass it to the prompt. To ensure that modifications were applied minimally, i.e. the difference in performance can be attributed to a particular linguistic feature change rather than to a large number of changes, we calculate Levenshtein distance between original and modified samples, discarding those which changed more than expected. Finally, we employ thorough human checks to ensure that modifications are valid and meaningful. For classification tasks we ran this validation in two stages on Prolific [^7] https://www.prolific.com/, with thorough quality control and four annotators per sample (see Section˜A.2 for annotation details, inter-annotator agreement and quality statistics). During the first stage, we check if resulting texts have the required modification and are still fluent, rejecting samples that do not meet these criteria. We find that most tests have a high retain rate of 70-90%. During the second stage, we ask annotators to select the tasks’ label (such as positive or negative for sentiment analysis) for modified samples, without knowing the labels for original samples, to check if the ground-truth label changed (e.g. after applying negation the sentiment can change from positive to negative). To make sure the task is solvable after modification, i.e. a label is not ambiguous or hard to choose, we add an option such as “Neutral” for sentiment analysis or “Not sure” for other tasks, and reject samples where annotators chose it.
For generation tasks, to demonstrate how the framework can be applied with minimal human intervention, we skip the first stage of human checks and only use tests with reliable modifications (over 70% retain rate). For the second stage, the authors of the paper validate modified samples in terms of whether the modification changes the label (for example, if it changes the premises in the math task, leading to a different answer) and whether the task is still solvable (for example, if there are modifications that prevent correct instruction following). Examples of such valid and rejected modifications, as well as modifications that led to label change, are shown in Table˜9 in Appendix. Such setup reduces the human effort, making the framework generalizable to new tasks, while still ensuring the validity of modifications.
@@ 4.2 Tested models and evaluation
We experiment with three smaller PLMs (BERT, GPT-2, and T5), three base LLMs (GPT-4o, Llama 3.1 and Claude 3.5), and two reasoning LLMs (GPT-5, DeepSeek R1). Both PLMs and LLMs are tested on classification tasks, while only LLMs are used for generation tasks. The pretrained models are fine-tuned on the respective training set with an additional task-specific classification layer on top (Section˜A.5 in Appendix). The LLMs are prompted with task-related instructions in a zero-shot way (Table˜10).
@@
We follow the original metrics used for the task to define correctness. For dialogue contradiction detection, coreference resolution, sentiment analysis, and GSM the correctness metric is accuracy. For NER, we use per-sample entity-level mean F1 (Nakayama, 2018), while IFEval employs strict success (all constraints satisfied) as correctness.
@@ 5.1 Use Case: NER and Coreference tasks
Table 2: NER: Unrobustness (U, %) by model and modification. Claude stands for Claude 3.5, Llama for Llama 3.1, and DS for DeepSeek R1. Full results with confidence interval (CI) in Table˜13.
@@
Table 3: Coreference: Unrobustness (U, %) by model and modification. Notation as in Table˜2. Full results with CI in Table˜14.
@@ 5.2 Use case: GSM and IFEval
We find that on the GSM task (Table˜4) models demonstrate major brittleness to Negation tasks, where all of them fail to take into account the changed or even reversed logic of the premise, showing memorisation effects. Base LLMs such as Claude-3.5 and especially Llama 3.1 are also affected by context changes, such as failing to accomodate for changed names and currencies in Geographical bias test, or explaining the problem in informal style or Singlish dialect. Interestingly, reasoning LLMs are more robust to such changes.
@@
Table 4: GSM: Unrobustness (U, %) by model and modification. Notation as in table˜2. Full results with CI in Table˜17.
@@
Table 5: IFEval: Unrobustness (U, %) by model and modification. Notation in Table˜2. Full results with CI in Table˜18.
@@ 6.1 Model brittleness is task-specific
We call our framework task-agnostic, since the same set of prompts can be used to generate a battery of tests applicable to all tasks. This, however, does not mean that all tests are meaningful for all the tasks. As the results above (and in Section˜A.8) show, the tests that actually “fired", or showed substantial drops or increases, are different for each task. Even test such as negation which is known to be difficult for models (Ravichander et al., 2022; Truong et al., 2022) has almost no effect on the NER task. On the other hand, seemingly “simple” tests (capitalization) that have a negligible effect on most of the tasks led to some dramatic drops on the NER and IFEval tasks. Thus, our framework is task-agnostic because it can be applied to reveal lack of robustness on a new task without having a preconceived (and potentially wrong) notion of what tests are meaningful for it.
@@ 6.2 LLMs are not always more robust than PLMs, and reasoning enhanced training does not always help
In classification tasks, where PLMs were finetuned on the original dataset, one would expect larger fluctuations in performance compared to LLMs, as modifications are likely to change the data distribution and break the identical distribution assumption that the fine-tuning process is based on. This holds for two out of four classification tasks we tested — Coreference resolution (Table˜3) and Sentiment analysis (Table˜16 in Section˜A.8) — where PLMs are overall less robust to perturbations. However, on two other classification tasks — NER (Table˜2) and Dialogue contradiction detection (Table˜15 in Section˜A.8) — while base LLMs are overall more robust than PLMs, reasoning LLMs are as brittle as the latter. Moreover, while LLMs are on average (across all models and tests) better (see the “Average” row in Table˜20), they are sometimes markedly brittle to a particular test. For example, Claude-3.5 is substantially more affected by concept modifications than any PLMs on Coreference resolution, or GPT-4o affected by grammatical role modifications on Dialogue contradiction task.
Comparing between base and reasoning LLMs across tasks does not reveal a clear winner either. Reasoning seems to improve overall robustness on tasks such as GSM (Table˜4), Coreference (Table˜3) or Sentiment Analysys (Table˜16); however, on other tasks the reasoning models perform similar to base LLMs (IFEval, Table˜5) or even worse than them (NER, Table˜2, and Dialogue contradiction detection, Table˜15).
We also test if a reasoning model (GPT-5) is able to handle the modifications better if it is aware of them. To do that, we include additional context into the prompt where we alert the model that the sample was modified, and explicitly specify the type of modification (e.g. by telling it that the active voice was changed to passive). We found no consistent gain when we do this (see Appendix Table˜12).
6.3 Scaling helps but with diminishing effects and only for superficial modifications
To test if increasing the parameter size helps with robustness, we additionally compare Llama-3.1 405B (which was used in all experiments) with two smaller versions – Llama-3.1-8B and Llama-3.1-70B – on the GSM task (the results in Table˜6). Overall, we observe a large improvement in robustness when going from 8B $\rightarrow$ 70B but observe diminishing effects when going beyond that. Low-level orthographic and syntactic modifications benefit strongly from increased model size. In contrast, more complex perturbations, especially negation, remain highly challenging even at 405B. This suggests that scaling primarily improves surface-level robustness, while deeper semantic changes remain brittle.
Category Modification Llama-8b Llama-70b Llama-405b Avg
Bias Temporal 8.0 5.0 4.0 5.7
Geographical 8.0 3.0 7.0 6.0
Length 8.0 2.0 2.0 4.0
Orthogr. Spelling 10.0 2.0 2.0 4.7
Capitalization 11.0 2.0 5.0 6.0
Punctuation 5.0 1.0 5.0 3.7
Syntax Conjunction 12.0 2.0 4.0 6.0
Voice 11.0 2.0 6.0 6.3
Semantic Concept 12.0 1.0 5.0 6.0
Negation 22.0 17.0 18.0 19.0
Discourse Appraisal 7.0 5.0 1.0 4.3
Varieties Style 16.0 6.0 6.0 9.3
Dialect 10.0 6.0 6.0 7.3
Average 10.8 4.2 5.5 6.8
Table 6: GSM Scaling: Unrobustness (U, %) by Llama model size and modification. Notation in Table˜2. Full results with CI in Table˜19.
@@ 6.4 Models are overall less robust to natural, linguistically valid modifications
Some of the work on robustness focuses on corrupting the text in adversarial way, such as shuffling characters or changing the case (Belinkov and Bisk, 2017; Li et al., 2018; Eger and Benz, 2020), while some includes natural, lingustically equivalent perturbations such as synonym replacement (Ren et al., 2019; Jin et al., 2020). Our framework combines both types of tests and allows to compare their efficiency for a particular task. We find that on average linguistically valid modifications reveal unrobustness more often than corrupting the text (i.e. such tests as Orthography) (see “Average” column in Table˜20). The exception here are the NER task where orthography is directly relevant to performance, and, surprisingly, IFEval, where the model often confused the instructions regarding capitalization with capitalization used in the instruction. Thus, the choice of the test type – adversarial corruptions or natural modifications – should also be motivated by the downstream task.
@@ 7 Conclusion
In this paper, we present FLUKE, a task-agnostic, linguistically-driven framework that introduces minimal modifications to existing datasets to create a benchmark for assessing robustness. We automate these modifications by prompting GPT-4o, followed by thorough manual validation to ensure data quality. To showcase the generalizability of this framework, we apply it to create modified instances across six tasks covering both classification and generation, and evaluate the performance of smaller fine-tuned models and large language models. Our findings reveal that: (1) the impact of modifications is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) while LLMs demonstrate better overall robustness compared to PLMs, they still exhibit significant brittleness, and on some tasks reasoning LLMs are less robust than base LLMs; (3) increasing the model size helps only with more surface-level modifications and has diminishing effects; (4) models are overall less robust to linguistically valid modifications than to adversarial corruption, but this again depends on the task; (5) the ability of a model to implement a linguistic feature during generation does not correlate with its robustness to that feature on downstream tasks. We hope that FLUKE provides an alternative approach for evaluating models, and that model developers would consider integrating these results into model cards (Mitchell et al., 2019) when releasing models.